OpenStudy (anonymous):

Find the quotient of -40x^3-12x^2+16x/-4x

OpenStudy (anonymous):

A. 10x^2 + 3x − 4 B. −40x^3 − 12x^2 − 4 C. −10x^2 + 3x − 4 D. 10x^2 − 3x + 4

OpenStudy (harsimran_hs4):

you can cancel out x from numerator and denominator and denominator and then divide numerator by -4(this is what is left in the denominator after x gets canceled out)

OpenStudy (harsimran_hs4):

can you proceed now?

OpenStudy (anonymous):

I don't really know what the 'numerator' and 'denominator' are.

OpenStudy (harsimran_hs4):

oops that is what i did not expect it`s taught way back in 2 or 3rd grade or even earlier e.g \[\frac{ a }{ b }\] here a is numerator and b is denominator

OpenStudy (anonymous):

I know haha I remember learning about it I just hadn't heard it be used since then so I didn't remember. Now how do you cancel out x?

OpenStudy (harsimran_hs4):

-40x^3-12x^2+16x/-4x = -40x^2-12x+16/-4 = 10x^2 + 3x -4 clear??

OpenStudy (anonymous):

yes, thank you!

OpenStudy (harsimran_hs4):

:)